# The drastic impact of Eddington-limit induced mass ejections on massive star populations

### v1

@@ 2.3 Model variations
Throughout this paper, we investigate how different choices of the input physics affect the synthetic stellar populations of the SMC and LMC. To do so, we create the following sets of stellar evolution models:
  • Model Eject: This is our standard model, which uses the input physics as described in the previous sections, including the modeling of Eddington-limit induced mass ejections. We created an additional synthetic population for the SMC called SMC-SFH, where we assume that star formation ended $3.5\,\mathrm{Myr}$ ago to match the observed population better.
  • Model no-Eject: Same setup as detailed in Sect. 2.1, but without the Eddington-limit induced mass ejections.
  • Model Binary: A combination of our Model Eject with the LMC binary grids from mar2:16 and pau1:22. For more details, refer to Sect. 2.4.2.
@@
  • Model sc10: Same setup as in Model Eject, but we changed the efficiency of semiconvective mixing to $\alpha_{\mathrm{sc}}=10$.
@@ 2.4.1 Single stars
To test our physical setup, we create synthetic populations of single stars per initial metallicity and per model variation, using the different single-star grid of detailed MESA models (see Sects. 2.1 and 2.3). The resulting synthetic populations are compared to the observed populations of the LMC and SMC. For our each of our single-star populations, each of the 33 detailed single-star models of a grid is assigned a statistical weight based on its formation probability and the duration of each time step. The probability of forming a star with a specific mass is determined by the initial mass function (IMF, sal1:55). In this work, we adopt a kro1:01 IMF and, for simplicity, a constant star formation rate (SFR) of $\mathrm{SFR}_{\mathrm{LMC}}=0.1\,{\,M_{\odot}\,\mbox{yr}^{-1}}$ for the LMC (har1:09, corrected for a Kroupa IMF) and $\mathrm{SFR}_{\mathrm{SMC}}=0.05\,{\,M_{\odot}\,\mbox{yr}^{-1}}$} for the SMC (rub1:18). The time spent in a specific evolutionary phase is determined by the detailed evolutionary models and does not rely on any external assumptions. More details on how the weights for individual models are calculated can be found in Appendix B.1.
@@ 2.4.2 Binary stars
Most massive stars are born in binary or higher-order multiple systems (san2:12; san1:13; san1:14; san1:25; dun1:15; fro1:25; vil1:25). If they have sufficiently close separations, the components interact during their lifetimes. Calculating a large grid of binary models spanning the full initial period and initial mass ratio parameter space is beyond the scope of this paper. However, to test the impact of binarity, we combine our single-star grids with already calculated MESA binary evolutionary grids at LMC metallicity. For the initial primary mass range from $M_{\mathrm{1,\,ini}}\approx 10\,M_{\odot}\text{\,--\,}25\,M_{\odot}$, we employ the binary evolutionary grid presented in mar2:16, and for the range $M_{\mathrm{1,\,ini}}\approx 28\,M_{\odot}\text{\,--\,}89\,M_{\odot}$, we use the grid from pau1:22, which reproduces the observed luminosity distribution of the LMC WR population. These grids have the same mass spacing ($\Delta\log(M_{\mathrm{ini}}/M_{\odot})=0.05$) as our single-star grids. The models from mar2:16 span an initial period range from $\log(P_{\mathrm{orb,\,ini}}/\mathrm{d})=0.15\,\text{--}\,3.5$, while the more massive star models from pau1:22 that can evolve to larger stellar radii span initial periods from $\log(P_{\mathrm{orb,\,ini}}/\mathrm{d})=0.15\,\text{--}\,4.0$. Both grids cover the full initial period space where binary interactions occur. We use a period spacing of $\Delta\log(P_{\mathrm{orb,\,ini}}/\mathrm{d})=0.05$. Furthermore, both grids cover initial mass ratios in the range from $q_{\mathrm{ini}}=0.25\,\text{--}\,0.95$, with a spacing of $\Delta q_{\mathrm{ini}}=0.05$. In total, these binary grids encompass about 18 000 detailed MESA models. To model mass-transfer, both binary model grids employ the “contact” scheme of MESA, allowing, as the name suggests, for contact phases. Furthermore, it is assumed that the accretion of mass is rotationally limited, leading to quasi-conservative mass-transfer in short-period systems and non-conservative mass-transfer in wide-period systems. Note that these grids use some different physical assumptions compared to our single-star models, such as mass-loss rates, and, thus, detailed comparisons should be made with caution.
@@ 2.5 Definitions of distinct stellar groups
For a consistent comparison of observations to the synthetic populations, we apply selection criteria to both populations to differentiate between the various evolutionary phases. The evolutionary phases considered in this work include WR stars, along with their multiple subtypes, as well as O-type stars, very massive stars (VMS), and supergiants that are blue (BSG), yellow (YSG), and red (RSG). Figure 2 provides an overview of the regions occupied by the different stellar groups. A detailed explanation on how stars are assigned to each group is provided in Appendix C.
@@ 3.1 The LMC population
If not mentioned otherwise, the works used to calculate the numbers of observed stars are the same as the ones quoted in Fig. 1. (b) Stars with spectral type Of/WN were excluded from this number, as these are not covered by our $\tau$-criterion (see Appendix C). (c) Only 7 of the WC/WO stars have been analyzed with non-LTE stellar atmosphere models, but 26 are confirmed (bre1:81; tes1:93; bar1:82; neu1:12). (d) dor1:13 investigated the completeness of the O-star content in the 30 Dor region of the LMC and reported the presence of about $N_{\mathrm{O,\,30Dor}}=570$ O stars. Furthermore, we know that 30 Dor hosts $N_{\mathrm{WR,\,30Dor}}=25$ WN/WC/WO stars (dor1:13, and references therein). Assuming that WR stars (excluding the extremely young Of/WN stars) are tracers of massive star populations, we infer that the number of O-stars in the LMC can be approximated as $N_{\mathrm{O,\,LMC}}=N_{\mathrm{WR,\,LMC}}/N_{\mathrm{WR,\,30Dor}}\cdot N_{\mathrm{O,\,30Dor}}$. The number quoted here corresponds to the average of this range. (e) The BLOeM sample, i.e., the largest spectroscopic sample of massive stars in the SMC, contains 159 O-stars and has an estimated completeness of around $\sim 35\%$ compared to Gaia DR3. This implies that there are $\sim 500$ O-type stars in the SMC (she1:24). Note that this number might be larger since Gaia cannot resolve the dense cores of young clusters.
@@
Figure 3: HRDs of the massive star population (symbols) in the LMC (left), and SMC (right), compared to the synthetic population of the Model Eject (contours). A selection of the same stellar evolution models is shown as solid lines, color-coded by their surface abundances. Small black dots on the evolution tracks mark equidistant timesteps of $\Delta t=$30\,000\text{\,}\mathrm{y}\mathrm{r}$\$. The hatched region indicates the empirical HD limit (hum1:79). The different symbols mark different stellar groups and have the same sources as mentioned in Fig. 1. In the SMC’s population, we added the WO star DR1 from IC 1613 (tra1:13), a galaxy which has SMC-like metallicity and only this WR star.
@@
Figure 4: Empirical luminosity distributions for the different cool supergiants (diamonds) in the LMC (left) and SMC (right) are compared to the predictions from the Model Eject (hatched bins). The error margins give the 84.13% confidence upper and lower margins of a Poissonian distribution accounting for small numbers (isr1:68). The sources of the observed stars are the same as in Fig. 1.
@@
Deficiencies of our single-star population are the lack of OB supergiants near the terminal age main-sequence (TAMS), as well as the BSG and YSG populations, when compared to the observed ones. san1:13 reported a lack of binaries in the O9.7 I, suggesting that they could be post-interaction binaries or merger products. Alternatively, the BSG might be explainable when using different assumptions on mixing (gil1:21). We discuss this in more detail in Sect. 4 and Appendix D. At the high-mass end, the predicted number of 28 VMS stars differs by a factor of two from the observed value of 13. At such high masses, the population is very sensitive to the assumed form of the IMF and star formation history (SFH). Small changes in these assumptions can easily account for the difference of the factor of two.
The synthetic single-star population of the LMC contains about 132 WR stars, which is close to the observed number of 122 WRs. The different WR subtypes serve as an excellent probe to test the stellar physics used in models, as they are sensitive to the assumptions made about mass-loss and mixing. Figure 5 shows the luminosity distributions of the WR subtypes in the LMC and SMC. Our population covers roughly the luminosity distribution of the observed WR stars and can explain the so far unexplained faintest single WR stars (she1:20, and references therein). There are only small discrepancies: a few too many H-poor WN stars around $\log(L/L_{\odot})\approx 6.1$, originating from the core hydrogen-burning WN population after the Eddington-limit induced mass ejections, and a few too bright H-free WNs around $\log(L/L_{\odot})\approx 5.8$ originating from stars experiencing mass ejections. Both of these features, as well as the overabundance of VMS in the synthetic population compared to observations, might suggest that our assumptions on the IMF or SFH are too simplistic, leading to such artifacts. The low number of predicted fainter H-free WN stars in the LMC population may be related to our choice of when to transition from the hot star mass-loss recipe to the WR mass-loss recipe. However, shifting this transition to lower $\Gamma_{\text{\!e}}$ values would lead to an overproduction of WN stars in the SMC, which is inconsistent with observations. Another crucial factor is the impact of binary interactions, which can lead to a more efficient removal of the H-rich envelope, making it easier for the WN stars to become H-free. pau1:22 report that binary interactions could be the primary mechanism behind the WN and WC populations in the LMC. We discuss this further in Sect. 4.2. For the WC/WO population in the LMC, our models predict that only stars that undergo a mass ejection episode evolve into WC/WO-type stars. Note that the brightest WC/WO stars in the LMC are most likely binaries, leading to too high luminosities when applying the bolometric correction (bar1:01; bar2:01; pau1:22).
@@ 3.2 The SMC population
Figure 6: HRD of the SMC population (left) and luminosity distributions of the WR subtypes in the SMC (right) compared to the model predictions of our Model Eject, where star-formation stopped $3.5\,\mathrm{Myr}$ ago. In the SMC’s WC/WO population, we added the WO star DR1 from IC 1613 (tra1:13). The symbols and colors are the same as in Figs. 3 and 5.
@@
As in the LMC, models with $M_{\mathrm{ini}}\gtrsim 50\,M_{\odot}$ experience a mass ejection episode, preventing most stars from exceeding the HD limit. Envelope inflation is $Z$-dependent, causing stars to expand less rapidly at lower metallicity. This can be seen by a few models that surpass the HD limit for less than a few thousand years. In total, one star is predicted beyond the HD limit, consistent with observations. Given the weak $Z$ dependence of the Eddington-limit induced mass ejection prescription, it might still be possible that below SMC metallicity stars can be found beyond the HD limit. However, this number should still be much smaller compared to models that ignore Eddington-limit induced mass ejections (see Sect. 4.1). Furthermore, the synthetic SMC population predicts 17 VMS, whereas none have been observed. Although we are working within a parameter regime where low-number statistics are relevant, the absence of any observed stars that should be the brightest in the galaxy is concerning.
@@ 3.3 The SMC population with changed SFH
The overabundance of O-type stars, VMS, and WR stars, each arising from stars with $M_{\mathrm{ini}}\gtrsim 28\,M_{\odot}$, in the synthetic SMC population, may suggest that our assumption of a constant SFH could be erroneous. To explore this further, we created a new synthetic population, assuming that star formation in the SMC stopped $3.5\,\mathrm{Myr}$ ago. We chose this stopping condition as it matches well the shape of the observed stars in the SMC and is roughly consistent with the age estimates of the bright O-stars in the young cluster NGC 346 (ric1:23). The HRD and WR luminosity distribution for this new population are shown in Fig. 6.
@@
Figure 7: HRDs of the massive star population (symbols) in the LMC (left), and SMC (right), compared to the synthetic population of the Model no-Eject (contours). The symbols and colors have the same meaning as in Fig. 3.
From the HRD of the synthetic population, illustrated in the left panel of Fig. 6, it is immediately apparent that the age cut removes all VMS stars (their evolutionary lifetime is much shorter than our age cut), bringing the synthetic population into better agreement with observations. Additionally, the region near the ZAMS, for stars with $M_{\mathrm{ini}}\gtrsim 25\,M_{\odot}$, is no longer populated, and the contours of the synthetic population now match the observed distribution much more closely. Our synthetic population now contains about 890 O-type stars, which is closer to the estimated number of 500 O-type stars, but still not in full agreement. Upon closer inspection of the HRD shown in Fig. 6, an absence of O-stars in the spectroscopic sample can be seen near the predicted turn-off point located around $\log(T/\mathrm{kK})\approx 4.5$ and $\log(L/L_{\odot})\approx 5.7$. However, this absence is not present in the photometric Gaia sample of sch1:21, suggesting it might be a coincidence that these stars are missing in our spectroscopic sample. Alternatively, this lack of stars could reflect a binarity effect, where the most luminous stars are binary products, namely post-interaction binaries or mergers. If it were a binary effect, it could also imply that our SFH is still too active and thus might lead to an overestimation of the O-type stars.
@@
Figure 8: Empirical luminosity distributions for the different cool supergiants (diamonds) in the LMC (left) and SMC (right) are compared to the predictions from the models with different physical setups (bins). The sources of the observed stars are the same as in Fig. 1.

### v2

@@ 2.3 Model variations
Throughout this paper, we investigate how different choices of the input physics affect the synthetic stellar populations of the SMC and LMC. To do so, we created the following sets of stellar evolution models:
  • Model Eject: This is our standard model. It uses the input physics as described in the previous sections, including the modeling of Eddington-limit induced mass ejections. These models assume a constant star-formation history (SFH). We created an additional synthetic population for the SMC called SMC-SFH, where we assume that star formation ended $3.5\,\mathrm{Myr}$ ago to match the observed population better.
  • Model no-Eject: This model has the same setup as detailed in Sect. 2.1, but without the Eddington-limit induced mass ejections.
  • Model Binary: This model is a combination of our Model Eject with the LMC binary grids from mar2:16 and pau1:22. We provide more details in Sect. 2.4.2.
@@
  • Model sc10: This model has the same setup as in Model Eject, but we changed the efficiency of semiconvective mixing to $\alpha_{\mathrm{sc}}=10$.
@@ 2.4.1 Single stars
To test our physical setup, we create synthetic populations of single stars per initial metallicity and per model variation, using the different single-star grid of detailed MESA models (see Sects. 2.1 and 2.3). The resulting synthetic populations are compared to the observed populations of the LMC and SMC. For our each of our single-star populations, each of the 33 detailed single-star models of a grid is assigned a statistical weight based on its formation probability and the duration of each time step. The probability of forming a star with a specific mass is determined by the initial mass function (IMF; sal1:55). In this work, we adopt a kro1:01 IMF and, for simplicity, a constant star formation rate (SFR) of $\mathrm{SFR}_{\mathrm{LMC}}=0.1\,{\,M_{\odot}\,\mbox{yr}^{-1}}$ for the LMC (har1:09, corrected for a Kroupa IMF) and $\mathrm{SFR}_{\mathrm{SMC}}=0.05\,{\,M_{\odot}\,\mbox{yr}^{-1}}$} for the SMC (rub1:18). The time spent in a specific evolutionary phase is determined by the detailed evolutionary models and does not rely on any external assumptions. More details on how the weights for individual models are calculated can be found in Appendix B.1.
@@ 2.4.2 Binary stars
Most massive stars are born in binary or higher-order multiple systems (san2:12; san1:13; san1:14; san1:25; dun1:15; fro1:25; vil1:25). If they have sufficiently close separations, the components interact during their lifetimes. Calculating a large grid of binary models spanning the full initial period and initial mass ratio parameter space is beyond the scope of this paper. However, to test the impact of binarity, we combine our single-star grids with already calculated MESA binary evolutionary grids at LMC metallicity. For the initial primary mass range from $M_{\mathrm{1,\,ini}}\approx 10\,M_{\odot}\text{\,--\,}25\,M_{\odot}$, we employ the binary evolutionary grid presented in mar2:16, and for the range $M_{\mathrm{1,\,ini}}\approx 28\,M_{\odot}\text{\,--\,}89\,M_{\odot}$, we use the grid from pau1:22, which reproduces the observed luminosity distribution of the LMC WR population. These grids have the same mass spacing ($\Delta\log(M_{\mathrm{ini}}/M_{\odot})=0.05$) as our single-star grids. The models from mar2:16 span an initial period range from $\log(P_{\mathrm{orb,\,ini}}/\mathrm{d})=0.15\,\text{--}\,3.5$, while the more massive star models from pau1:22 that can evolve to larger stellar radii span initial periods from $\log(P_{\mathrm{orb,\,ini}}/\mathrm{d})=0.15\,\text{--}\,4.0$. Both grids cover the full initial period space where binary interactions occur. We use a period spacing of $\Delta\log(P_{\mathrm{orb,\,ini}}/\mathrm{d})=0.05$. Furthermore, both grids cover initial mass ratios in the range from $q_{\mathrm{ini}}=0.25\,\text{--}\,0.95$, with a spacing of $\Delta q_{\mathrm{ini}}=0.05$. In total, these binary grids encompass about 18 000 detailed MESA models. To model mass-transfer, both binary model grids employ the “contact” scheme of MESA, allowing, as the name suggests, for contact phases. Furthermore, it is assumed that the accretion of mass is rotationally limited, leading to quasi-conservative mass-transfer in short-period systems and non-conservative mass-transfer in wide-period systems. Note that these grids use some different physical assumptions compared to our single-star models, such as mass-loss rates, and thus, detailed comparisons should be made with caution.
@@ 2.5 Definitions of distinct stellar groups
For a consistent comparison of observations to the synthetic populations, we apply selection criteria to both populations to differentiate between the various evolutionary phases. The evolutionary phases considered in this work include WR stars, along with their multiple subtypes, as well as O-type stars, very massive stars (VMSs), and supergiants that are blue (BSG), yellow (YSG), and red (RSG). Figure 2 provides an overview of the regions occupied by the different stellar groups. A detailed explanation on how stars are assigned to each group is provided in Appendix C.
@@ 3.1 The LMC population
If not mentioned otherwise, the works used to calculate the numbers of observed stars are the same as the ones quoted in Fig. 1. (b) Stars with spectral type Of/WN were excluded from this number, as these are not covered by our $\tau$-criterion (see Appendix C). (c) Only 7 of the WC/WO stars have been analyzed with non-local thermodynamic equilibrium stellar atmosphere models, but 26 are confirmed (bre1:81; tes1:93; bar1:82; neu1:12). (d) dor1:13 investigated the completeness of the O-star content in the 30 Dor region of the LMC and reported the presence of about $N_{\mathrm{O,\,30Dor}}=570$ O stars. Furthermore, we know that 30 Dor hosts $N_{\mathrm{WR,\,30Dor}}=25$ WN/WC/WO stars (dor1:13, and references therein). Assuming that WR stars (excluding the extremely young Of/WN stars) are tracers of massive star populations, we infer that the number of O-stars in the LMC can be approximated as $N_{\mathrm{O,\,LMC}}=N_{\mathrm{WR,\,LMC}}/N_{\mathrm{WR,\,30Dor}}\cdot N_{\mathrm{O,\,30Dor}}$. The number quoted here corresponds to the average of this range. (e) The BLOeM sample, i.e., the largest spectroscopic sample of massive stars in the SMC, contains 159 O-stars and has an estimated completeness of around $\sim 35\%$ compared to Gaia DR3. This implies that there are $\sim 500$ O-type stars in the SMC (she1:24). Note that this number might be larger since Gaia cannot resolve the dense cores of young clusters.
@@
Figure 3: Hertzsprung-Russell diagrams of the massive star population (symbols) in the LMC (left) and SMC (right) compared to the synthetic population of the Model Eject (contours). A selection of the same stellar evolution models is shown as solid lines, color coded by their surface abundances. Small black dots on the evolution tracks mark equidistant timesteps of $\Delta t=$30\,000\text{\,}\mathrm{y}\mathrm{r}$\$. The hatched region indicates the empirical HD limit (hum1:79). The different symbols mark different stellar groups and have the same sources as mentioned in Fig. 1. In the SMC’s population, we added the WO star DR1 from IC 1613 (tra1:13), a galaxy that has an SMC-like metallicity and only this WR star.
@@
Figure 4: Empirical luminosity distributions for the different cool supergiants (diamonds) in the LMC (left) and SMC (right) compared to the predictions from the Model Eject (hatched bins). The error margins represent the upper and lower bounds of the 84.13% confidence interval of a Poisson distribution, accounting for small numbers (isr1:68). The sources of the observed stars are the same as in Fig. 1.
@@
Deficiencies of our single-star population are the lack of OB supergiants near the terminal age main-sequence as well as the BSG and YSG populations when compared to the observed ones. san1:13 reported a lack of binaries in the O9.7 I, suggesting that they could be post-interaction binaries or merger products. Alternatively, the BSG might be explainable when using different assumptions on mixing (gil1:21). We discuss this in more detail in Sect. 4 and Appendix D. At the high-mass end, the predicted number of 28 VMSs differs by a factor of two from the observed value of 13. At such high masses, the population is very sensitive to the assumed form of the IMF and SFH. Small changes in these assumptions can easily account for the difference of the factor of two.
The synthetic single-star population of the LMC contains about 132 WR stars, which is close to the observed number of 122 WRs. The different WR subtypes serve as an excellent probe to test the stellar physics used in models, as they are sensitive to the assumptions made about mass-loss and mixing. Figure 5 shows the luminosity distributions of the WR subtypes in the LMC and SMC. Our population covers roughly the luminosity distribution of the observed WR stars and can explain the so far unexplained faintest single WR stars (she1:20, and references therein). There are only small discrepancies: a few too many H-poor WN stars around $\log(L/L_{\odot})\approx 6.1$, which originate from the core hydrogen-burning WN population after the Eddington-limit induced mass ejections, and too few bright H-free WNs around $\log(L/L_{\odot})\approx 5.8$, which originate from stars experiencing mass ejections. Both of these features, as well as the overabundance of VMSs in the synthetic population compared to observations, might suggest that our assumptions on the IMF or SFH are too simplistic and thus lead to such artifacts. The low number of predicted fainter H-free WN stars in the LMC population may be related to our choice of when to transition from the hot star mass-loss recipe to the WR mass-loss recipe. However, shifting this transition to lower $\Gamma_{\text{\!e}}$ values would lead to an overproduction of WN stars in the SMC, which is inconsistent with observations. Another crucial factor is the impact of binary interactions, which can lead to a more efficient removal of the H-rich envelope, making it easier for the WN stars to become H-free. pau1:22 report that binary interactions could be the primary mechanism behind the WN and WC populations in the LMC. We discuss this further in Sect. 4.2. For the WC/WO population in the LMC, our models predict that only stars that undergo a mass ejection episode evolve into WC/WO-type stars. Note that the brightest WC/WO stars in the LMC are most likely binaries, leading to too high luminosities when applying the bolometric correction (bar1:01; bar2:01; pau1:22).
@@ 3.2 The SMC population
Figure 6: Hertzsprung-Russell diagram of the SMC population (left) and luminosity distributions of the WR subtypes in the SMC (right) compared to the model predictions of our Model Eject, where star-formation stopped $3.5\,\mathrm{Myr}$ ago. In the SMC’s WC/WO population, we added the WO star DR1 from IC 1613 (tra1:13). The symbols and colors are the same as in Figs. 3 and 5.
@@
As in the LMC, models with $M_{\mathrm{ini}}\gtrsim 50\,M_{\odot}$ experience a mass ejection episode, preventing most stars from exceeding the HD limit. Envelope inflation is $Z$-dependent, causing stars to expand less rapidly at lower metallicity. This can be seen by a few models that surpass the HD limit for less than a few thousand years. In total, one star is predicted beyond the HD limit, consistent with observations. Given the weak $Z$ dependence of the Eddington-limit induced mass ejection prescription, it might still be possible that below SMC metallicity stars can be found beyond the HD limit. However, this number should still be much smaller compared to models that ignore Eddington-limit induced mass ejections (see Sect. 4.1). Furthermore, the synthetic SMC population predicts 17 VMSs, whereas none have been observed. Although we are working within a parameter regime where low-number statistics are relevant, the absence of any observed stars that should be the brightest in the galaxy is concerning.
@@ 3.3 The SMC population with changed SFH
The overabundance of O-type stars, VMSs, and WR stars, each arising from stars with $M_{\mathrm{ini}}\gtrsim 28\,M_{\odot}$, in the synthetic SMC population, may suggest that our assumption of a constant SFH could be erroneous. To explore this further, we created a new synthetic population, assuming that star formation in the SMC stopped $3.5\,\mathrm{Myr}$ ago. We chose this stopping condition as it matches well the shape of the observed stars in the SMC and is roughly consistent with the age estimates of the bright O-stars in the young cluster NGC 346 (ric1:23). The HRD and WR luminosity distribution for this new population are shown in Fig. 6.
@@
Figure 7: Hertzsprung-Russell diagrams of the massive star population (symbols) in the LMC (left) and SMC (right) compared to the synthetic population of the Model no-Eject (contours). The symbols and colors are the same as in Fig. 3.
From the HRD of the synthetic population, illustrated in the left panel of Fig. 6, it is immediately apparent that the age cut removes all VMSs (their evolutionary lifetime is much shorter than our age cut), bringing the synthetic population into better agreement with observations. Additionally, the region near the ZAMS, for stars with $M_{\mathrm{ini}}\gtrsim 25\,M_{\odot}$, is no longer populated, and the contours of the synthetic population now match the observed distribution much more closely. Our synthetic population now contains about 890 O-type stars, which is closer to the estimated number of 500 O-type stars, but still not in full agreement. Upon closer inspection of the HRD shown in Fig. 6, an absence of O-stars in the spectroscopic sample can be seen near the predicted turn-off point located around $\log(T/\mathrm{kK})\approx 4.5$ and $\log(L/L_{\odot})\approx 5.7$. However, this absence is not present in the photometric Gaia sample of sch1:21, suggesting it might be a coincidence that these stars are missing in our spectroscopic sample. Alternatively, this lack of stars could reflect a binarity effect, where the most luminous stars are binary products, namely post-interaction binaries or mergers. If it were a binary effect, it could also imply that our SFH is still too active and thus might lead to an overestimation of the O-type stars.
@@
Figure 8: Empirical luminosity distributions for the different cool supergiants (diamonds) in the LMC (left) and SMC (right) compared to the predictions from the models with different physical setups (bins). The sources of the observed stars are the same as in Fig. 1.